OpenStudy (my4babiesandme):

f(x)+ 10 - x^2 find the vertex? line of semetry? and min/max?

OpenStudy (anonymous):

is this \[f(x)=10-x^2\]?

OpenStudy (my4babiesandme):

yes

OpenStudy (anonymous):

I think you mean f(x)=-10 x^2 , if so then -1/10 y = x^2 ==> 4a=1/10 ==> a=1/40 ===> vertex =(0,0) , line of symmtry is x=0 , maximum is (0,0)

OpenStudy (anonymous):

this is a parabola that faces down. looks just like \[y=x^2\] but flipped over and raised ten units. vertex is (0, 1) and the axis of symmetry is x = 0 aka the y- axis

OpenStudy (anonymous):

there is no minimum value but the maximum is 10 when x = 0

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D10-x^2